OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = x2 – 2x + 1, what is f(g(x)) ? 4x2 – 8x + 5 4x2 + 2x + 2 x2 + 2x + 2 x2 – 2x + 2

OpenStudy (anonymous):

4(x^2-2x+1)+1=f(g(x)) 4x^2-8x+4+1=f(g(x)) f(g(x))=4x^2-8x+5

OpenStudy (anonymous):

Sub g(x) into wherever x is inside f(x)

OpenStudy (anonymous):

That makes no sense to me lol is it the first answer?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

let's call g(x) something you are more familiar with, like g(x) = y, then sub that into f(g(x)) Which gives you f(y)=4x+1

OpenStudy (anonymous):

Since you substitute whatever is inside f() into any x you have in your function, you'll get 4(y)+1=f(y) Then just resub y=g(x) = x^2 – 2x + 1